OpenStudy (poopsiedoodle):

How might one go about doing this problem? Solve for x: x^2 + 24x + 90 = 0

OpenStudy (anonymous):

Factor first :)

OpenStudy (anonymous):

What factors of 90, when added together, give you 24?

OpenStudy (poopsiedoodle):

Gimme a second.

OpenStudy (anonymous):

Take your time (:

OpenStudy (poopsiedoodle):

I don't think any do. 18 and 5 is the closest you can get to 24, which is 23.

OpenStudy (anonymous):

Hmm, hold on a sec

OpenStudy (poopsiedoodle):

I will do that.

OpenStudy (anonymous):

Well, since it isn't factorable, we have to use the quadratic formula. Have you learned it?

OpenStudy (poopsiedoodle):

Hardly. I've just started on it.

OpenStudy (anonymous):

Well, this may take some explaining to do, so hold on for a bit haha.

OpenStudy (anonymous):

x^2 + 24x + 90 = 0 Let's break down what this problem means. By plugging in a certain number(s) for x, you get zero. It's pretty much like plotting a graph of x^2 + 24x + 90 and looking for where the graph touches zero. Did I lose you yet?

OpenStudy (anonymous):



OpenStudy (poopsiedoodle):

Okay. I get it so far.

OpenStudy (anonymous):

Alright. So since we have an equation that isn't factorable, we find the zeroes by plugging it into the quadratic formula.

OpenStudy (anonymous):

This is the quadratic formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac} }{ 2a }\] Now, a, b, and c are the numbers in your equation -- \[a^2+bx+c\] In this case: \[x^2 + 24x + 90\] so, a=1, b= 24, and c=90.

OpenStudy (poopsiedoodle):

Ok, lemme try it now.

OpenStudy (anonymous):

Remember, the ± symbol indicates that there are two expressions! (ex. 1 ± 3 is 1 + 3 AND 1 - 3)

OpenStudy (poopsiedoodle):

\[\Large0= \frac{ -24 \pm \sqrt{24^{2} -4(90)} }{ 2 }\]

OpenStudy (anonymous):

Good so far!

OpenStudy (anonymous):

except, it's x =

OpenStudy (poopsiedoodle):

Well, since we're already using x in the original problem, how about Y?

OpenStudy (anonymous):

But we're trying to solve for x, haha

OpenStudy (poopsiedoodle):

eh, whatever. By the way, I'm having a slight problem with \[\sqrt{4(90)}\]

OpenStudy (anonymous):

That's just \[\sqrt{4 • 90} = \sqrt{360}\]

OpenStudy (poopsiedoodle):

Ah. \[\Large x= \frac{ -24 \pm 24\sqrt{360} }{ 2 }\]

OpenStudy (anonymous):

No, remember 24^2 is under the radical (Don't take it out!) \[\sqrt{24^2 - 360}\]